# Beyond Freezing: Sparse Tuning Enhances Plasticity in Continual Learning with Pre-Trained Models

### v1

@@ 4.3 Plugging MIST into PTM-based Continual Learning: The Algorithm
5 Experiment
5.1 Experimental Setups
Benchmark. We consider five representative benchmark datasets and randomly split each of them into 10 disjoint tasks. Specifically, CIFAR-100 dataset Krizhevsky and Hinton (2009) consists of 100-class natural images with 500 training samples per class. ImageNet-R dataset Hendrycks et al. (2021a) contains 200-class images, spliting into 24,000 and 6,000 images for training and testing, respectively. ImageNet-A Hendrycks et al. (2021b) dataset consists of 200 classes and contains 7,500 adversarially filtered images, which are known to significantly degrade the performance of machine learning models. CUB-200 Wah et al. (2011) dataset includes 200-class bird images with around 60 images per class, 30 of which are used for training and the rest for testing. Cars-196 Krause et al. (2013) dataset includes 196 types of car images, split into 8,144 and 8,040 images for training and testing, respectively. Performance is evaluated using the standard CL metric, Average Accuracy Chaudhry et al. (2019), defined as: $A_{t}=\frac{1}{t}\sum_{i=1}^{t}R_{t,i}$, where $R_{t,i}$ denotes the classification accuracy on the $i$-th task after training on the $t$-th task. We report both $A_{T}$ and $\bar{A}$ in the main paper. Here, $\bar{A}$ denotes the mean of $A_{t}$ over all tasks: $\bar{A}=\frac{1}{T}\sum_{t=1}^{T}A_{t}$. It reflects the average accuracy of all classes seen so far after each incremental task.
Implementation. Following previous works Wang et al. (2022b, a), we adopt a pre-trained ViT-B/16 backbone Dosovitskiy et al. (2020) for all baselines. For continual learning on downstream tasks, we follow the original implementations by employing the Adam optimizer for L2P, DualPrompt, and CoDA-Prompt, and the SGD optimizer for all other baselines. Our method, MIST, is inserted as a plug-in module before each selected baseline and is trained for 20 epochs using the SGD optimizer with a learning rate of 0.0001. In the MI-based selection stage, we select the top $k\%=5\%$ most sensitive parameters. For each mini-batch, we further apply a dropout rate of $d\%=90\%$ to the selected parameters, resulting in only 0.5% of total parameters being updated per batch. MIST solely optimizes the MI loss (Eq. 8), with the temperature $\tau$ set to 0.5. For each task, MIST first conducts this sparse fine-tuning, after which the corresponding baseline resumes training using its original configuration. We adopt this setting consistently across all datasets in our experiments.
@@ 5.1 Experimental Setups
$\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$
CODA-Prompt Smith et al. (2023) 91.3 86.9 78.5 73.4 63.9 52.7 84.1 79.3 52.1 45.4
SLCA++ Zhang et al. (2024a) 94.1 91.5 83.0 77.5 67.1 58.7 91.0 86.7 79.2 73.8
APER(Adapter) Zhou et al. (2025) 83.9 85.9 74.2 66.9 62.4 52.1 90.5 85.6 52.8 40.5
L2P Wang et al. (2022b) 86.7 83.3 74.5 68.6 53.9 44.9 81.7 67.4 53.9 39.6
@@
SimpleCIL Zhou et al. (2025) 87.1 81.3 61.1 54.3 59.8 48.5 90.9 85.6 38.8 27.8
@@
5.2 Experimental Results
Overall performance. To assess the versatility of MIST, we plug it into five representative freeze-based methods: L2P Wang et al. (2022b), DualPrompt Wang et al. (2022a), SLCA Zhang et al. (2023a), RanPAC McDonnell et al. (2023), and SimpleCIL Zhou et al. (2025). Among them, L2P and DualPrompt are prompt-based methods that freeze the PTM and learn token-like prompts for adaptation. SimpleCIL does not involve any parameter tuning and directly trains a prototype classifier on frozen representations. RanPAC is an adapter-based method that inserts and fine-tunes lightweight modules in the PTM. SLCA adopts a full fine-tuning strategy with a reduced learning rate to balance stability and plasticity. As shown in Table 1, incorporating MIST consistently improves all methods across all datasets. For example, DualPrompt/MIST achieves accuracy gains of +1.6%, +6.0%, +4.4%, +1.4%, and +11.2% on CIFAR100, ImageNet-R, ImageNet-A, CUB200, and Cars196, respectively. Furthermore, we observe that many methods perform poorly on the Cars196 dataset. For instance, the final accuracies $A_{T}$ of L2P and SimpleCIL are only 39.6% and 27.8%, respectively. This is mainly because pretraining knowledge offers limited utility for complex fine-grained vehicle classification, making it particularly challenging for models to adapt to such new domains. After inserting MIST as a pre-adaptation stage, the final accuracies of L2P and SimpleCIL increase by +13.1% and +15.7% respectively, indicating that MIST effectively enhances the model’s ability to align with domain-specific structures before classifier training. Among all methods, RanPAC/MIST achieves the best overall performance, indicating that even well-designed adapter-based methods benefit from the MI-guided tuning stage. This highlights the complementary nature of MIST as a general plug-in for enhancing adaptation in various PTM-based CL frameworks.
@@ 5.2 Experimental Results
Effect of MIST To better understand the effect of MIST on PTM-based CL performance, we visualize both the new task accuracy and the incremental accuracy after each task in Figure 4 and Figure 4. As shown, MIST consistently improves the learning effectiveness across all inserted methods. Specifically, in Figure 4, the new task accuracy increases significantly for all methods after integrating MIST, demonstrating its ability to enhance the model’s adaptability to newly arrived tasks. This improvement indicates that the pre-adaptation phase provided by MIST helps the PTM align more effectively with task-specific distributions. Correspondingly, Figure 4 shows that MIST also leads to notable gains in incremental accuracy across all tasks. This is attributed to the improved learning efficiency on new tasks, which in turn contributes to higher cumulative accuracy when evaluated on all seen classes. Taken together, these results highlight the effectiveness of MIST as a plug-in component that improves the task-specific learning capacity of freeze-based methods while preserving their stability, ultimately leading to consistent performance gains in CL scenarios.
Table 2: Comparison of different tuning strategies used as pre-adaptation before RanPAC. FFT fine-tunes all parameters, while other methods update only the 5% parameters per task.
Method CIFAR100 ImageNet-R ImageNet-A CUB200 Cars196
$\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$ $\bar{A}$ $A_{T}$
RanPAC 94.0 90.8 83.2 77.9 70.1 61.4 92.6 88.9 82.8 74.6
+FFT 61.0 39.6 57.6 36.7 32.0 11.3 51.9 28.6 31.3 11.2
+Grad 56.2 34.0 54.8 33.9 43.6 28.9 50.8 26.9 29.9 10.6
+Rand 45.3 20.9 52.4 11.8 34.8 13.0 68.1 34.9 33.7 10.9
+L2 42.0 15.9 39.0 14.4 23.1 8.2 61.6 38.0 29.6 10.4
+MIST 95.3 92.4 84.9 81.0 72.5 62.5 93.6 90.4 83.0 76.4
@@
Comparison with naive sparse tuning strategies We integrate different pre-adaptation strategies into RanPAC, including full fine-tuning (FFT) of all parameters, top 5% selection based on gradient magnitude (Grad) or parameter norm (L2), and random 5% selection (Rand), and evaluate their performance across multiple datasets, and evaluate their performance across multiple datasets. As shown in Table 2, all alternative methods perform significantly worse than MIST, underperforming the baseline without any pre-adaptation. MIST consistently outperforms these strategies because it leverages MI to assess parameter sensitivity, thereby achieving effective task adaptation while minimizing disruptions to the pre-trained representations. In contrast, these alternative methods do not consider preserving the pre-trained knowledge of the PTM when selecting parameters, making them prone to catastrophic forgetting and performance collapse. Overall, MIST offers a more balanced adaptation path by jointly preserving plasticity and stability, demonstrating its superiority as a general plug-in pre-adaptation module.
Table 3: Ablation studies on Imagenet-R.
MI Sparse MI loss Dropout $A_{T}$
77.9
66.8
76.7
81.0
@@
Ablation study. Table  4 present the ablation study. When applying only the MI sparse selection, the model achieves 66.8% accuracy. Although sparse selection reduces parameter interference, the optimization remains guided by the cross-entropy loss, which as discussed in Eq.(5), fails to explicitly preserve the pre-trained feature distribution. Nevertheless, this approach still outperforms alternative selection strategies, as evidenced in Table 2. Upon introducing the MI loss, performance improves to 76.7%, indicating that the MI objective effectively guides the model toward downstream distributions while retaining generalization. Finally, incorporating gradient dropout further improves the accuracy to 81.0%, as it regularizes the update path and mitigates overfitting to static parameter importance. These findings confirm that each component of MIST—MI-guided sparsity, MI loss, and dropout—contributes synergistically to overall performance improvements.
Parameter efficiency analysis. Table 4 compares the efficiency of different methods in terms of (1) $\Delta$P: the number of parameters updated per mini-batch (in millions), (2) FLOPs: flops for updating the selected parameters per batch, and (3) Time: time required to train a batch on an NVIDIA RTX 4090 GPU. Among the methods, SLCA performs full fine-tuning and thus has the highest update cost—both in terms of parameters (85.40M) and time (12.3ms). L2P only updates prompt tokens, but incurs additional overhead (12.4ms) likely due to its key-query matching mechanism during token routing. MIST, while updating only 0.43M parameters per task, incurs slightly higher computation time (9.7ms) compared to RanPAC. This is because computing the MI loss requires augmented views of each sample. In summary, MIST achieves the lowest update cost without introducing any additional parameters, making it easily pluggable into other methods.
@@
Effect of selection rate and dropout rate. We conduct a hyperparameter study to explore how the parameter selection rate $k\%$ and the gradient dropout rate $d\%$ affect the performance of MIST, as reported in Table 6 and Table 6, respectively. We observe that high sparsity generally yields better performance. For instance, selecting only 5% of parameters per task achieves $A_{T}=81.0\%$ on ImageNet-R and $76.4\%$ on Cars196, which outperforms full fine-tuning. This validates that MIST is able to effectively adapt to new tasks while preserving the pre-trained structure by updating only a small subset of critical parameters. The performance of $k=5\%$ is very close to $k=10\%$, with only marginal differences. Given that fewer parameters are involved and the computational cost is lower, we adopt $k=5\%$ as the default in practice. Table 6 shows that increasing the gradient dropout rate significantly improves performance, especially from $d=0\%$ to $d=90\%$. This confirms that dropout acts as an effective regularizer, helping to suppress local gradient bias and mitigate overfitting to static parameter importance scores. The best performance is observed when $d=90\%$, and we use this setting as the default for all experiments.
6 Conclusion
In this paper, we investigate the fundamental challenge of balancing plasticity and generalization in PTM-based CL. We reveal that direct fine-tuning often compromises the pre-trained feature distribution, while existing freeze-based methods suffer from limited adaptability to new tasks. Through a theoretical lens grounded in MI, we analyze how gradients derived from MI objectives offer a more stable optimization path by avoiding unnecessary perturbations to the PTM. Motivated by this, we propose Mutual Information-guided Sparse Tuning, a lightweight and plug-and-play pre-adaptation strategy that selectively updates only the most informative parameters before each incremental task. By computing an MI-based Fisher Information Matrix, MIST identifies sensitive parameters, then applies strong gradient dropout to regularize the update path, enabling the PTM to better align with task-specific distributions while maintaining generalizable representations. Extensive experiments demonstrate that MIST can be seamlessly integrated into various freeze-based CL frameworks, consistently boosting performance across diverse datasets, especially under large distribution shifts. Moreover, MIST achieves this without introducing any additional parameters, and with minimal computational cost, making it highly efficient and practical. The limitation of MIST lies in its reliance on efficient approximation of the Fisher matrix. When the data within a task exhibits significant distributional variation, this approximation may become inaccurate, potentially compromising the effectiveness of MIST. In the future, we plan to explore more robust Fisher estimation techniques that can adapt to intra-task variation.
Appendices
Appendix A Gradient of Mutual Information
$I(X;Y)=\sum_{x,y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)},$ (A1)
$\frac{\partial I(X;Y)}{\partial\theta}$ $=\sum_{x,y}\frac{\partial p(x,y)}{\partial\theta}\log\frac{p(x,y)}{p(x)p(y)}+\sum_{x,y}p(x,y)\cdot\frac{\partial}{\partial\theta}\log\frac{p(x,y)}{p(x)p(y)}$ (A2)
Note that:
$\sum_{x,y}p(x,y)\cdot\frac{1}{p(x)}\cdot\frac{\partial p(x)}{\partial\theta}$ $=\sum_{x}\left(\frac{\partial p(x)}{\partial\theta}\cdot\sum_{y}\frac{p(x,y)}{p(x)}\right)$ (A3)
Similarly,
$\sum_{x,y}p(x,y)\cdot\frac{1}{p(y)}\cdot\frac{\partial p(y)}{\partial\theta}=\sum_{y}\frac{\partial p(y)}{\partial\theta}\cdot\sum_{x}\frac{p(x,y)}{p(y)}=\sum_{y}\frac{\partial p(y)}{\partial\theta}.$ (A4)
$\sum_{x}\frac{\partial p(x)}{\partial\theta}=0,\quad\sum_{y}\frac{\partial p(y)}{\partial\theta}=0.$ (A5)
The last two terms in Eq. A2 cancel out, and the gradient simplifies to:
$\frac{\partial I(X;Y)}{\partial\theta}=\sum_{x,y}\frac{\partial p(x,y)}{\partial\theta}\log\frac{p(x,y)}{p(x)p(y)}.$ (A6)
$F_{\mathrm{MI}}=\left(\mathbb{E}_{x\sim\mathcal{D}_{t}}[g(x)]\right)^{2}.$ (A7)
$F^{\prime}_{\mathrm{MI}}=\left(\frac{1}{N}\sum_{j=1}^{N}g(x_{j})\right)^{2}.$ (A8)
$\mathbb{E}[F^{\prime}_{\mathrm{MI}}]$ $=\mathbb{E}\left[\left(\frac{1}{N}\sum_{j=1}^{N}g(x_{j})\right)^{2}\right]$ (A9)
Appendix C Common tuning strategies for PTMs
Table A1: Final accuracy achieved under different batch sizes
batch size 4 32 64
Final accuracy (%) 42.0 43.1 43.5
MI can only be estimated at the batch level. The supervised InfoNCE loss defined in Eq. (8) estimates MI $I(X;Y)$ using only mini-batch samples, limiting its representation of the true data distribution. Prior studies Guo et al. [2022], Oord et al. [2018] and our theoretical result in Eq. (A6) indicate that more diverse batches improve MI estimation quality, as empirically confirmed in Table A1, where increasing the batch size from 4 to 64 raises final accuracy from 42.0% to 43.5%. Nevertheless, batch-wise computation inherently involves the term $\partial p(x;\theta)/\partial\theta^{i}$, causing unavoidable perturbations to the pretrained feature structure.
Sparse tuning helps mitigate the above limitations. To address these issues, we adopt MI-guided parameter selection and gradient dropout as regularization strategies. Specifically, we select only the top-$k\%$ most MI-sensitive parameters and randomly drop $d\%$ of them in each batch, resulting in only $0.5\%$ of parameters being updated. This strong sparsity reduces the risk of feature drift, mitigates overfitting to local updates, and preserves structural integrity during task adaptation. Hence, sparse tuning not only improves efficiency but also plays a crucial role in stabilizing the adaptation process, making it an essential complement to batch-level MI optimization. Figure A1 presents the final accuracy comparisons on ImageNet-R and Cars196. Across all baseline methods, incorporating MIST consistently yields notable accuracy enhancements. This highlights the general effectiveness of our proposed MIST strategy in improving incremental learning performance.
(a) ImageNet-R dataset
(b) Cars dataset
Figure A1: Performance comparison of incremental learning methods with and without the proposed MIST. The inclusion of MIST consistently improves final accuracy across all evaluated methods and datasets.
Appendix E Experiments on Domain-Incremental Learning

### v2

@@ Introduction
Experiment
Experimental Setups
Benchmark. We consider five representative benchmark datasets and randomly split each of them into 10 disjoint tasks. Including CIFAR-100 (krizhevsky2009cifar), ImageNet-R (hendrycks2021many), ImageNet-A (hendrycks2021natural), CUB-200 (wah2011cub) and Cars-196 (krause2013cars). Performance is evaluated using the standard CL metric, Average Accuracy (shim2021ASER), defined as: $A_{t}=\frac{1}{t}\sum_{i=1}^{t}R_{t,i}$, where $R_{t,i}$ denotes the classification accuracy on the $i$-th task after training on the $t$-th task. We report both $A_{T}$ and $\bar{A}$ in the main paper. Here, $\bar{A}$ denotes the mean of $A_{t}$ over all tasks: $\bar{A}=\frac{1}{T}\sum_{t=1}^{T}A_{t}$. It reflects the average accuracy of all classes seen so far after each incremental task.
Implementation. Following previous works (wang2022l2p; wang2022dualprompt), we adopt a pre-trained ViT-B/16 backbone (dosovitskiy2020vit) for all baselines. Except for OCM (guo2022ocm), no methods use a buffer; OCM is implemented in an offline setting with a buffer size of 1000. We follow the original implementations by employing the Adam optimizer for L2P, DualPrompt, and CoDA-Prompt, and the SGD optimizer for all other baselines. Our method, MIST, is inserted as a plug-in module before each selected baseline and is trained for 20 epochs using the SGD optimizer with a learning rate of 0.0001. In the MI-based selection stage, we select the top $k\%=5\%$ most sensitive parameters. For each mini-batch, we further apply a dropout rate of $d\%=90\%$ to the selected parameters, resulting in only 0.5% of total parameters being updated per batch. MIST solely optimizes the MI loss (Eq. 7), with the temperature $\tau$ set to 0.5. For each task, MIST first conducts this sparse fine-tuning, after which the corresponding baseline resumes training using its original configuration. We adopt this setting consistently across all datasets in our experiments.
@@
$\bar{A}$ $_{T} A¯\bar{A}$ $_{T} A¯\bar{A}$ $_{T} A¯\bar{A}$ $_{T} A¯\bar{A}$ $A_{T}$
EWC (ewc_0) 53.2 41.2 42.3 26.2 22.0 7.9 57.6 37.8 46.7 24.3
OCM (guo2022ocm) 63.6 37.0 77.0 66.7 58.1 44.4 85.1 71.7 65.3 50.5
CODA-Prompt (smith2023coda-prompt) 91.3 86.9 78.5 73.4 63.9 52.7 84.1 79.3 52.1 45.4
SLCA++ (zhang2024slca++) 94.1 91.5 83.0 77.5 67.1 58.7 91.0 86.7 79.2 73.8
APER(Adapter) (zhou2025revisiting) 83.9 85.9 74.2 66.9 62.4 52.1 90.5 85.6 52.8 40.5
L2P (wang2022l2p) 86.7 83.3 74.5 68.6 53.9 44.9 81.7 67.4 53.9 39.6
@@
EASE (ease) 91.8 87.4 81.1 75.1 65.3 54.9 90.9 85.8 38.4 27.3
+MIST 92.2 88.2 81.8 76.3 66.0 56.9 91.8 87.0 57.0 47.2
SimpleCIL (zhou2025revisiting) 87.1 81.3 61.1 54.3 59.8 48.5 90.9 85.6 38.8 27.8
@@
Experimental Results
Overall performance. To assess the versatility of MIST, we plug it into six representative freeze-based methods: L2P, DualPrompt, SLCA, EASE, RanPAC, and SimpleCIL. Among them, L2P and DualPrompt are prompt-based methods that freeze the PTM and learn token-like prompts for adaptation. SimpleCIL does not involve any parameter tuning and directly trains a prototype classifier on frozen representations. RanPAC and EASE are adapter-based methods that inserts and fine-tunes lightweight modules in the PTM. SLCA adopts a full fine-tuning strategy with a reduced learning rate. As shown in Table 1, incorporating MIST consistently improves all methods across all datasets. For example, DualPrompt/MIST achieves accuracy gains of +1.6%, +6.0%, +4.4%, +1.4%, and +11.2% on the five datasets, respectively. Furthermore, we observe that many methods perform poorly on the Cars196 dataset. For instance, the $A_{T}$ of L2P and SimpleCIL are only 39.6% and 27.8%, respectively. This is mainly because pretraining knowledge offers limited utility for complex fine-grained vehicle classification, making it particularly challenging for models to adapt to such new domains. After inserting MIST, the final accuracies of L2P and SimpleCIL increase by +13.1% and +15.7% respectively, indicating that MIST effectively enhances the model’s ability to align with domain-specific structures before classifier training. Among all methods, RanPAC/MIST achieves the best overall performance, indicating that even well-designed adapter-based methods benefit from the MI-guided tuning stage.
@@
Effect of MIST. To better understand the effect of MIST on PTM-based CL performance, we visualize both the new task accuracy (Figure 4) and the incremental accuracy (Figure 4). New task accuracy refers to the accuracy on the newly learned task, while incremental accuracy denotes the average accuracy over all tasks learned so far. As shown, MIST improves the learning effectiveness across all inserted methods. Specifically, in Figure 4, the new task accuracy increases significantly for all methods after integrating MIST. This improvement indicates that the pre-adaptation phase provided by MIST helps the PTM align more effectively with task-specific distributions. Correspondingly, Figure 4 shows that MIST also leads to notable gains in incremental accuracy across all tasks. This is attributed to the improved learning efficiency on new tasks, which in turn contributes to higher cumulative accuracy when evaluated on all seen classes. Taken together, these results highlight the effectiveness of MIST as a plug-in component that improves the task-specific learning capacity of freeze-based methods while preserving their stability, ultimately leading to consistent performance gains in CL scenarios.
Table 2: Comparison of tuning strategies used for pre-adaptation in RanPAC. “MS” refers to MI-based sparse selection, “ML” to MI loss, and “Drop” to dropout. “MS+ML+Drop” corresponds to our MIST implementation. (a) shows naive selection strategies, (b) adds MI loss and dropout to a naive strategy, and (c) uses MI-based sparse selection combined with MI loss and dropout.
Method ImageNet-R Cars196
$\bar{A}$ $_{T} A¯\bar{A}$ $A_{T}$
RanPAC 83.2 77.9 82.8 74.6
+FFT 57.6 36.7 31.3 11.2
(a) +Rand 52.4 11.8 33.7 10.9
+L2 39.0 14.4 29.6 10.4
+Grad 54.8 33.9 29.9 10.6
(b) +Grad+ML 76.7 67.1 64.2 40.3
+Grad+ML+Drop 77.5 68.3 64.9 42.0
+MS 75.5 66.8 76.0 66.2
(c) +MS+ML 82.2 76.7 81.3 72.1
+MS+ML+Drop 84.9 81.0 83.0 76.4
@@
Comparison with naive sparse tuning strategies. We integrate several pre-adaptation strategies into RanPAC, including full fine-tuning (FFT) of all parameters, top 5% selection based on gradient magnitude (Grad) or parameter norm (L2), and random 5% selection (Rand). As shown in Table 2(a) and (b), all these alternatives perform significantly worse than MIST. This is because MIST leverages mutual information (MI) to assess parameter sensitivity, enabling effective task adaptation while minimizing disruptions to the pre-trained representations. In contrast, the alternative methods ignore the importance of preserving pre-trained knowledge during parameter selection, making them vulnerable to catastrophic forgetting and performance collapse. Overall, MIST offers a more balanced adaptation path by jointly preserving plasticity and stability, demonstrating its superiority as a general plug-in pre-adaptation module.
Ablation study. Table 2(c) presents the ablation study of MIST. When applying only MI-based sparse selection (”MS”), the model achieves the lowest accuracy among the three. Although sparse selection reduces parameter interference, the optimization remains guided by the cross-entropy loss, which fails to explicitly preserve the pre-trained feature distribution. Nevertheless, this approach still outperforms alternative selection strategies. Upon introducing the MI loss (”MS+ML”), performance improves, indicating that the MI objective effectively guides the model toward downstream distributions while retaining generalization. Finally, adding gradient dropout (”MS+ML+Drop”) yields the highest accuracy across all settings, as it regularizes the update path and mitigates overfitting to static parameter importance. To further validate the effectiveness of MI-based sparse selection, we also apply MI loss and dropout to the Grad-based strategy in Table 2(b). Although this leads to some performance gains, its results remain clearly inferior to MIST. This highlights the unique advantage of MI-guided selection and the complementary role of the three components in MIST. These findings confirm that MI sparsity, MI loss, and dropout contribute synergistically to overall performance gains.
@@
[…]
Conclusion
In this paper, we investigate the fundamental challenge of balancing plasticity and generalization in PTM-based CL. We reveal that direct fine-tuning often compromises the pre-trained feature distribution, while existing freeze-based methods suffer from limited adaptability to new tasks. Through a theoretical lens grounded in MI, we analyze how gradients derived from MI objectives offer a more stable optimization path by avoiding unnecessary perturbations to the PTM. Motivated by this, we propose Mutual Information-guided Sparse Tuning, a lightweight and plug-and-play pre-adaptation strategy that selectively updates only the most informative parameters before each incremental task. By computing an MI-based Fisher Information Matrix, MIST identifies sensitive parameters, then applies strong gradient dropout to regularize the update path, enabling the PTM to better align with task-specific distributions while maintaining generalizable representations. Extensive experiments demonstrate that MIST can be seamlessly integrated into various freeze-based CL frameworks, consistently boosting performance across diverse datasets, especially under large distribution shifts. The limitation of MIST lies in its reliance on efficient approximation of the Fisher matrix. When the data within a task exhibits significant distributional variation, this approximation may become inaccurate, potentially compromising the effectiveness of MIST. In the future, we plan to explore more robust Fisher estimation techniques that can adapt to intra-task variation.
Acknowledgment
This paper was supported by the National Science Foundation of China (No. 62406323), China Postdoctoral Science Foundation (No. 2024M753496), and Postdoctoral Fellowship Program of CPSF (No. GZC20232993).
Appendices
Appendix A Gradient of Mutual Information
Mutual information gradients help pre-trained models acquire new knowledge while resisting the collapse of pre-trained knowledge. Below, we provide a detailed mathematical derivation to support this claim.
$I(X;Y)=\sum_{x,y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)},$ (8)
$\frac{\partial I(X;Y)}{\partial\theta}$ $=\sum_{x,y}\frac{\partial p(x,y)}{\partial\theta}\log\frac{p(x,y)}{p(x)p(y)}+\sum_{x,y}p(x,y)\cdot\frac{\partial}{\partial\theta}\log\frac{p(x,y)}{p(x)p(y)}$ (9)
Note that:
$\sum_{x,y}p(x,y)\cdot\frac{1}{p(x)}\cdot\frac{\partial p(x)}{\partial\theta}$ $=\sum_{x}\left(\frac{\partial p(x)}{\partial\theta}\cdot\sum_{y}\frac{p(x,y)}{p(x)}\right)$ (10)
Similarly,
$\sum_{x,y}p(x,y)\cdot\frac{1}{p(y)}\cdot\frac{\partial p(y)}{\partial\theta}=\sum_{y}\frac{\partial p(y)}{\partial\theta}\cdot\sum_{x}\frac{p(x,y)}{p(y)}=\sum_{y}\frac{\partial p(y)}{\partial\theta}.$ (11)
$\sum_{x}\frac{\partial p(x)}{\partial\theta}=0,\quad\sum_{y}\frac{\partial p(y)}{\partial\theta}=0.$ (12)
The last two terms in Eq. 9 cancel out, and the gradient simplifies to:
$\frac{\partial I(X;Y)}{\partial\theta}=\sum_{x,y}\frac{\partial p(x,y)}{\partial\theta}\log\frac{p(x,y)}{p(x)p(y)}.$ (13)
$F_{\mathrm{MI}}=\left(\mathbb{E}_{x\sim\mathcal{D}_{t}}[g(x)]\right)^{2}.$ (14)
$F^{\prime}_{\mathrm{MI}}=\left(\frac{1}{N}\sum_{j=1}^{N}g(x_{j})\right)^{2}.$ (15)
$\mathbb{E}[F^{\prime}_{\mathrm{MI}}]$ $=\mathbb{E}\left[\left(\frac{1}{N}\sum_{j=1}^{N}g(x_{j})\right)^{2}\right]$ (16)
Appendix C Common tuning strategies for PTMs
Table 4: Final accuracy achieved under different batch sizes
batch size 4 32 64
Final accuracy (%) 42.0 43.1 43.5
MI can only be estimated at the batch level. The supervised InfoNCE loss defined in Eq. (8) estimates MI $I(X;Y)$ using only mini-batch samples, limiting its representation of the true data distribution. Prior studies (guo2022ocm; oord2018representation) and our theoretical result in Eq. (13) indicate that more diverse batches improve MI estimation quality, as empirically confirmed in Table 4, where increasing the batch size from 4 to 64 raises final accuracy from 42.0% to 43.5%. Nevertheless, batch-wise computation inherently involves the term $\partial p(x;\theta)/\partial\theta^{i}$, causing unavoidable perturbations to the pretrained feature structure.
Sparse tuning helps mitigate the above limitations. To address these issues, we adopt MI-guided parameter selection and gradient dropout as regularization strategies. Specifically, we select only the top-$k\%$ most MI-sensitive parameters and randomly drop $d\%$ of them in each batch, resulting in only $0.5\%$ of parameters being updated. This strong sparsity reduces the risk of feature drift, mitigates overfitting to local updates, and preserves structural integrity during task adaptation. Hence, sparse tuning not only improves efficiency but also plays a crucial role in stabilizing the adaptation process, making it an essential complement to batch-level MI optimization. Figure 5 presents the final accuracy comparisons on ImageNet-R and Cars196. Across all baseline methods, incorporating MIST consistently yields notable accuracy enhancements. This highlights the general effectiveness of our proposed MIST strategy in improving incremental learning performance.
(a) ImageNet-R dataset
(b) Cars dataset
Figure 5: Performance comparison of incremental learning methods with and without the proposed MIST. The inclusion of MIST consistently improves final accuracy across all evaluated methods and datasets.
@@ Appendix D Limitations of Batch-level MI Estimation and the Role of Sparse Tuning
Appendix E Effect of selection rate and dropout rate.
@@
We conduct a hyperparameter study to explore how the parameter selection rate $k\%$ and the gradient dropout rate $d\%$ affect the performance of MIST, as reported in Table 6 and Table 6, respectively. We observe that high sparsity generally yields better performance. For instance, selecting only 5% of parameters per task achieves $A_{T}=81.0\%$ on ImageNet-R and $76.4\%$ on Cars196, which outperforms full fine-tuning. This validates that MIST is able to effectively adapt to new tasks while preserving the pre-trained structure by updating only a small subset of critical parameters. The performance of $k=5\%$ is very close to $k=10\%$, with only marginal differences. Given that fewer parameters are involved and the computational cost is lower, we adopt $k=5\%$ as the default in practice. Table 6 shows that increasing the gradient dropout rate significantly improves performance, especially from $d=0\%$ to $d=90\%$. This confirms that dropout acts as an effective regularizer, helping to suppress local gradient bias and mitigate overfitting to static parameter importance scores. The best performance is observed when $d=90\%$, and we use this setting as the default for all experiments.
(a) ImageNet-A
(b) ImageNet-R
(c) Cifar-100
(d) Cub
Figure 6: Comparison of average forgetting F and accuracy A across datasets. F/M and A/M represent results with MIST integration. “Dual” denotes DualPrompt; “Ran” denotes RanPac.
Appendix F Impact of MIST on Forgetting and Learning Performance
To further compare the performance of different methods after integrating MIST, we report the average forgetting (kang2022forget) in Figure 6. Average Forgetting measures the degree of knowledge retention loss for each task after training on the final task, and is defined as:
$\text{Average Forgetting}=\frac{1}{T-1}\sum_{j=1}^{T-1}d_{T,j},$ (17)
where $d_{T,j}$ denotes the forgetting on task $j$ after training up to task $T$, computed as:
$d_{i,j}=\max_{k\in\{1,\ldots,i-1\}}a_{k,j}-a_{i,j},$ (18)
with $a_{i,j}$ representing the test accuracy on task $j$ after training on task $i$. This metric captures the worst-case drop in performance on previous tasks over the training trajectory.
As shown in the results, while incorporating MIST generally reduces the forgetting rate across most methods and datasets, there are instances where forgetting increases. This can be attributed to MIST’s core mechanism—enhancing the efficiency of learning new knowledge. In some cases, a lower forgetting rate without MIST does not imply better retention of prior knowledge. Instead, it may result from poor learning of new tasks, where limited improvement on recent tasks leads to a small performance gap between old and new tasks, thus producing an artificially low forgetting rate. Conversely, with MIST, some methods exhibit higher forgetting because MIST significantly boosts performance on new tasks. This amplifies the gap between old tasks (with stable or slightly degraded performance) and new tasks (with much improved accuracy), leading to a higher measured forgetting rate. Importantly, even in these cases, the overall average accuracy across all tasks improves with MIST. This reflects a more favorable balance between stability (retaining old knowledge) and plasticity (learning new knowledge). In incremental learning, a moderate increase in forgetting is acceptable—and often expected—when overall learning improves. It signals better adaptation to new tasks, which is the core goal: to acquire new knowledge effectively while maintaining reasonable retention of previous knowledge, rather than minimizing forgetting at the cost of stagnating new task performance.
Appendix G Experiments on Domain-Incremental Learning